OpenStudy (anonymous):

Find the calories in 16 potato chips if 10 potato chips contain 115 calories.

OpenStudy (luigi0210):

Well in one potato chip there is 11.5 calories. So 11.5*15=?

OpenStudy (phi):

I would use a ratio: \[ \frac{115 \text{ cals}}{10 \text{ chips }} = \frac{x \text{ cals}}{16 \text{ chips }} \]

OpenStudy (phi):

to solve, multiply both sides by 16 chips: \[ \frac{115 \text{ cals}}{10 \cancel{\text{ chips} }} \cdot 16 \cancel{\text{ chips} }= \frac{x \text{ cals}}{\cancel{16 \text{ chips }}} \cdot \cancel{16 \text{ chips }}\]

OpenStudy (phi):

you get \[ \frac{ 115 \cdot 16 }{10} \text{ cals } = x \]

OpenStudy (texaschic101):

I like proportions.... However, this is an indirect proportion..... 10/16 = 115/x cross multiply (10)(x) = (115)(16) 10x = 1840 x = 184 Lets check it : 10/16 = 115/184 cross multiply (10)(184) = (16)(115) 1840 = 1840 (correct) ANSWER: 16 potato chips contain 184 calories

OpenStudy (anonymous):

Thank you! :)

OpenStudy (texaschic101):

thats what we are here for :)